# Recalibrating binary probabilistic classifiers

### v2

@@ 4 Approaches to recalibration
Any solution as in (1b) to the recalibration problem together with the marginal feature distribution $Q(X)$ completely determines the target distribution $Q(X,Y)$. Since $Q(X,Y)\neq P(X,Y)$ in case $q\neq p$, any given solution specifies some distribution shift between the source and target distributions. Hence, in order to better understand the consequences of selecting a particular solution transformation $T$, it is natural to explore the solution approaches through the lens of distribution shift. In the following, we assume $10<\eta_{P}(x)<$ 1 for all $x\in\mathcal{X}$.
@@ 4.1 Recalibration under assumption of stretched or compressed covariate shift
At first glance recalibration of $\eta_{P}(X)$ to a fixed target prior probability $q$ might appear to be straightforward: Just define $\eta_{Q}(x)=\frac{q}{E_{Q}[\eta_{P}(X)]}\,\eta_{P}(x)$ for $x\in\mathcal{X}$, then $E_{Q}[\eta_{Q}(X)]=q$ immediately follows. This approach is called scaling (Section 3.1 of Ptak-Chmielewska and Kopciuszewski [15]).
@@
Recall that source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through covariate shift (in the sense of Storkey [21]) if it holds that $P[Y=y\,|\,X]=Q[Y=y\,|\,X]$ for all $y\in\mathcal{Y}$ with probability $1$ both under $P$ and $Q$. Hence if $t>1$ in (4a), one might call the implied distribution shift stretched covariate shift.
In case $t<1$ the induced distribution shift could be considered compressed covariate shift. In case $q>E_{Q}[\eta_{P}(X)]$, (4a) and (4b) imply $t>1$ and $\eta_{Q}(x)=1$ for all $x$ with $t\,\eta_{P}(x)\geq 1$. This may have the consequence of an unjustified increase of $AUC_{\eta_{Q}(X)}$ compared to $AUC_{\eta_{P}(X)}$ for the area under the curve AUC defined as in Section 3.3 below.
@@ 4.2 Recalibration under assumption of label shift
Source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through label shift (Lipton et al. [12]), previously called prior probability shift in the literature (Storkey [21]), if $P[X\in M\,|\,Y=y]=Q[X\in M\,|\,Y=y]$ for all measurable sets $M\subset\mathcal{X}$ and $y\in\mathcal{Y}$.
@@
for all measurable sets $M\subset\mathcal{X}$. If a prior probability $q=Q[Y=1]$ is given, then according to the posterior correction formula (Eq. (2.4) of Saerens et al. [18]), $\eta_{Q}$ is determined through (recall $p=P[Y=1]$)
@@
for all $x\in\mathcal{X}$ with probability $1$ under $Q$. In the credit risk community, the use of (6) for recalibration is popular (Section “Calibrating to PDs” of Bohn and Stein [3], Section 3.1 of Ptak-Chmielewska and Kopciuszewski [15]) because it avoids the problem of $\eta_{Q}(X)$ potentially taking the value $1$ for large $\eta_{P}(X)$ which may be encountered with capped scaling as in (4a). Cramer [6] (Sections 6.2 and 6.3) pointed out that in the context of logistic regression the related special case of (6) was known at least since 1979. Note that (6) implies (1b) with $T(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$.
@@
Proposition 1 is a consequence of (3a) in Section 3.3 as well as (1b) and (6). According to Proposition 1, recalibration under the assumption of label shift leaves the implied performance under the target distribution – sometimes called discriminatory power in the credit risk management literature (e.g. Bohn and Stein [3]) – unchanged when compared to the implied performance under the source distribution. In particular, Proposition 1 implies a necessary criterion for distribution shift to be label shift. Accordingly, in case $AUC_{S}\neq AUC_{S^{\ast}}$ source distribution $P$ and target distribution $Q$ cannot be related through label shift.
@@ 4.3 Recalibration under assumption of factorizable joint shift (FJS)
According to He at al. [10] and Tasche [25], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through factorizable joint shift (FJS) if there are functions $g:\mathcal{X}\to[0,\infty)$ and $b:\mathcal{Y}\to[0,\infty)$ such that $(x,y)\mapsto g(x)\,b(y)$ for $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ is a density of $Q(X,Y)$ with respect to $P(X,Y)$, i.e. it holds that $Q[(X,Y)\in M]=E_{P}\bigl{[}\mathbf{1}_{M}(X,Y)\,g(X)\,b(Y)\bigr{]}$ for all measurable sets $M\subset\mathcal{X}\times\mathcal{Y}$. See footnote 2 for the definition of the indicator $\mathbf{1}_{M}$.
By Corollary 4 of Tasche [25] and subject to mild technical conditions, under FJS the joint target distribution $Q(X,Y)$ is given by the target feature distribtuion $Q(X)$ and the class 1 posterior probability
@@
Note that recalibration of $\eta_{P}(X)$ to a target prior probability $q$ under the assumption of FJS works for arbritrary target feature distributions $Q(X)$. This is in stark contrast to recalibration under the assumption of label shift which only works when assuming that $Q(X)$ is given by (5). However, by Proposition 1 recalibration under the label shift assumption entails $AUC_{\eta_{Q}(X)}=AUC_{\eta_{P}(X)}$. The example in Section 5 below shows that AUC preservation is not in general true for recalibration under the FJS assumption.
(7a) implies (1b) with $T(\eta)=T_{\varrho}(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1}{\varrho}\,\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$. If $\varrho$ happens to take the value $1$, at first glance we are back in the context of label shift as in Section 4.2 above. However, as (5) need not hold true under the FJS assumption, in such cases there cannot be label shift. The type of shift modelled instead is called ‘invariant density ratio’ shift (Tasche [23]), defined as the special case of FJS with $\varrho=1$.
@@ 4.4 Recalibration under assumption of covariate shift with posterior drift (CSPD)
In Sections 4.1, 4.2 and 4.3, we identified the transformation $T$ of (1b) after we had described a recalibration method designed under assumption of one of three types of distribution shift, namely slightly modified covariate shift, label shift, and FJS. In contrast, in this section, we define $T$ and use it to characterise the type of distribution shift implied by the recalibration method.
According to Scott [19], the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through covariate shift with posterior drift (CSPD) if there is a strictly increasing transformation $T$ such that (1b) holds true.
As mentioned in Section 3.1, (1b) implies that $\eta_{P}(X)$ is sufficient for $X$ with respect to $Y$ under the target distribution $Q$. Since under CSPD the transformation $T$ is strictly increasing, (1b) also implies that $\eta_{Q}(X)$ and $\eta_{P}(X)$ are strongly comonotonic (Tasche [24]). As a consequence, Kendall’s $\tau$ and Spearman’s rank correlation both take the maximum value $1$ when applied to $(\eta_{Q}(X),\eta_{P}(X))$. This suggests that CSPD is a strong assumption which might be less often true than one would hope for. Nonetheless, comonotonicity of score and posterior probability is a common assumption in the literature. Chen et al. [5] call this assumption the rationality assumption.
If labels are available in the test dataset, under the CSPD assumption the transformation $T$ of (1b) can be approximately determined by means of isotonic regression. However, the general assumption for this paper is that there are no label observations for the instances in the test dataset. Therefore, we are going to apply a moment matching approach instead (quasi moment matching, QMM).
To be able to do so, we consider parametric CSPD where the transformation $T$ of (1b) is specified as follows through some strictly increasing and continuous distribution function $F$ on the real line and parameters $a,b\in\mathbb{R}$:
@@
This is not a radically new approach but rather a variation of what was called regression-based calibration by Ojeda et al. [14]. Here are two examples for natural choices of $F$ in (8a):
  • Logistic distribution function (inverse logit): $F(x)=\frac{1}{1+\exp(-x)}$, $x\in\mathbb{R}$. The parametric CSPD approach with inverse logit is often called ‘Platt scaling’ in the literature. However, as pointed out by Ojeda et al. [14], ‘Platt scaling’ sometimes also refers to the specification of $T$ as
@@
    In the following, we use the term Platt scaling to refer to (8b) and logistic CSPD to refer to (8a) with inverse logit.
  • Standard normal distribution function (inverse probit): $F(x)=\Phi(x)$, $x\in\mathbb{R}$. We refer to this choice of $F$ as normal CSPD.
@@
In (8c), AUC works like a second moment of the posterior probabilities. But in contrast to its effect on the second moment or the variance, the prior probability of the positive class has no effect on AUC. This makes the assumption of AUC invariance between source and target distributions more plausible (Tasche [22]).
Remark 1
@@
Then the mean of the resulting posterior probability $\eta_{Q}(X)=T_{a,0}(\eta_{P}(X))$ under the target distribution feature distribution $Q(X)$ is computed to obtain an estimate $\widehat{q}$ of $q$:
@@ 4.5 Quasi moment matching based on parametrised receiver operating characteristics
Tasche ([22], Section 5.2) modified an idea of van der Burgt [4] by assuming the ROC curve of a real-valued score $S$ to be
@@
for some fixed parameter $c\in\mathbb{R}$, with $\Phi$ denoting the standard normal distribution function. The ROC curve of (10a) emerges when the class-conditional score distributions in a binary classification problem are both univariate normal distributions with equal variances. But ROC curves like in (10a) may also be incurred in circumstances where the class-conditional distributions are not normal (Proposition 5.3 of Tasche [22]).
Tasche [22] showed that (10a) implies the following representation for the posterior probability given the score $S$ under the target distribution $Q$:
$Q[Y=1\,|\,S]\ =\ \frac{1}{1+\frac{1-q}{q}\,\exp\bigl{(}c^{2}/2-c\,\Phi^{-1}(F_{0}(S))\bigr{)}},$ (10b)
@@
Denote by $f$ an unconditional density of $S$ under $Q$ and by $f_{0}$ a density of $F_{0}$. Then the posterior probability $Q[Y=1\,|\,S=s]$ can be represented as
@@
In the case of scores $S$ under a discrete or empirical distribution, as described in Appendices 0.A.1 and 0.A.2, (11b) can be treated as a fixed point equation for the probabilities $Q[S=s\,|\,Y=0]=f_{0}(s)$ and be solved by a straightforward fixed-point iteration with intial values $Q[S=s]$, $s\in\mathcal{S}$. The numerical example of Section 5 below suggests that such an iterative approach converges as long as the prior probability $q$ is small and, hence, $f$ and $f_{0}$ are close to each other.
A further issue when making (10b) operational may occur when the class-conditional distribution function $F_{0}$ or any function approximating it takes the value $1$. In particular, this will happen if $F_{0}$ is approximated by an empirical distribution function. Then the term $\Phi^{-1}\bigl{(}F_{0}(s)\bigr{)}$ is ill-defined. See Appendix 0.A.2 for a workaround to deal with this issue.
Deploying a discrete distribution $F_{0}$ in (10b) as for instance an empirical approximation, makes it unlikely if not impossible to exactly match a pre-defined $AUC_{S}$ when using (10b). To control for the unavoidable deviation from the $AUC_{S}$ objective in this case we take recourse to the original QMM approach as proposed by Tasche [22]. Define
@@
To distinguish the two recalibration approaches presented in this section, we call the approach based on (10b) and (10c) ROC-based QMM and refer to the approach based on (12a) and (12b) as 2-parameter QMM.
5 Example
The example presented in this section illustrates the impact of the recalibration methods and assumptions discussed in Section 4 on the following characteristics of the target distribution:

### v3

@@ 4 Approaches to recalibration
Any solution as in (1b) to the recalibration problem together with the marginal feature distribution $Q(X)$ completely determines the target distribution $Q(X,Y)$. Since $Q(X,Y)\neq P(X,Y)$ in case $q\neq p$, any given solution specifies some distribution shift between the source and target distributions. Hence, in order to better understand the consequences of selecting a particular solution transformation $T$, it is natural to explore the solution approaches through the lens of distribution shift. In the following, the assumption $0<\eta_{P}(x)<1$ for all $x\in\mathcal{X}$ is made.
@@ 4.1 Recalibration under assumption of stretched or compressed covariate shift
At first glance recalibration of $\eta_{P}(X)$ to a fixed target prior probability $q$ might appear to be straightforward: Just define $\eta_{Q}(x)=\frac{q}{E_{Q}[\eta_{P}(X)]}\,\eta_{P}(x)$ for $x\in\mathcal{X}$, then $E_{Q}[\eta_{Q}(X)]=q$ immediately follows. This approach is called scaling (Section 3.1 of Ptak-Chmielewska and Kopciuszewski, 2022).
@@
Recall that source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through covariate shift (in the sense of Storkey, 2009) if it holds that P[Y=y|X]=Q[Y=y|X]P[Y=y\,|\,X]=Q[Y=y\,|\,X] for all $y\in\mathcal{Y}$ with probability 11 both under $P$ and $Q$. Hence if $t>1$ in (4a), one might call the implied distribution shift stretched covariate shift.
In case $t<1$ the induced distribution shift could be considered compressed covariate shift. In case $q>E_{Q}[\eta_{P}(X)]$, (4a) and (4b) imply $t>1$ and $\eta_{Q}(x)=1$ for all $x$ with $t\,\eta_{P}(x)\geq 1$. This may have the consequence of an undesired increase of $AUC_{\eta_{Q}(X)}$ compared to $AUC_{\eta_{P}(X)}$ for the area under the curve AUC defined as in Section 3.3 above.
@@ 4.2 Recalibration under assumption of label shift
Source distribution $P(X,Y)$ and target distribution $Q(X,Y)$ are related through label shift (Lipton et al., 2018), previously called prior probability shift in the literature (Storkey, 2009), if $P[X\in M\,|\,Y=y]=Q[X\in M\,|\,Y=y]$ for all measurable sets $M\subset\mathcal{X}$ and $y\in\mathcal{Y}$.
@@
for all measurable sets $M\subset\mathcal{X}$. If a prior probability $q=Q[Y=1]$ is given, then according to the posterior correction formula (Eq. (2.4) of Saerens et al., 2002), $\eta_{Q}$ is determined through (recall $p=P[Y=1]$)
@@
for all $x\in\mathcal{X}$ with probability $1$ under $Q$. In the credit risk community, the use of (6) for recalibration is popular (Section “Calibrating to PDs” of Bohn and Stein, 2009, Section 3.1 of Ptak-Chmielewska and Kopciuszewski, 2022) because it avoids the problem of $\eta_{Q}(X)$ potentially taking the value $1$ for large $\eta_{P}(X)$ which may be encountered with capped scaling as in (4a). Cramer (2003, Sections 6.2 and 6.3) pointed out that in the context of logistic regression the related special case of (6) was known at least since 1979. Note that (6) implies (1b) with $T(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$.
@@
Proposition 4.1 is a consequence of (3a) in Section 3.3 as well as (1b) and (6). According to Proposition 4.1, recalibration under the assumption of label shift leaves the implied performance under the target distribution – sometimes called discriminatory power in the credit risk management literature (e.g., Bohn and Stein, 2009) – unchanged when compared to the implied performance under the source distribution. In particular, Proposition 4.1 implies a necessary criterion for distribution shift to be label shift. Accordingly, in case $AUC_{S}\neq AUC_{S^{\ast}}$ source distribution $P$ and target distribution $Q$ cannot be related through label shift.
@@ 4.3 Recalibration under assumption of factorizable joint shift (FJS)
According to He et al. (2021) and Tasche (2022), the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through factorizable joint shift (FJS) if there are functions $g:\mathcal{X}\to[0,\infty)$ and $b:\mathcal{Y}\to[0,\infty)$ such that $(x,y)\mapsto g(x)\,b(y)$ for $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ is a density of $Q(X,Y)$ with respect to $P(X,Y)$, i.e. it holds that $Q[(X,Y)\in M]=E_{P}\bigl[\mathbf{1}_{M}(X,Y)\,g(X)\,b(Y)\bigr]$ for all measurable sets $M\subset\mathcal{X}\times\mathcal{Y}$. See footnote 2 for the definition of the indicator $\mathbf{1}_{M}$.
By Corollary 4 of Tasche (2022) and subject to mild technical conditions, under FJS the joint target distribution $Q(X,Y)$ is fully specified by the target feature distribtuion $Q(X)$ and the class 1 posterior probability
@@
Note that recalibration of $\eta_{P}(X)$ to a target prior probability $q$ under the assumption of FJS works for arbritrary target feature distributions $Q(X)$. This is in stark contrast to recalibration under the assumption of label shift which only works when assuming that $Q(X)$ is given by (5). However, by Proposition 4.1 recalibration under the label shift assumption entails $AUC_{\eta_{Q}(X)}=AUC_{\eta_{P}(X)}$. The example in Section 5 below shows that AUC preservation is not in general true for recalibration under the FJS assumption.
(7a) implies (1b) with $T(\eta)=T_{\varrho}(\eta)=\frac{\frac{p}{q}\,\eta}{\frac{p}{q}\,\eta+\frac{1}{\varrho}\,\frac{1-p}{1-q}\,(1-\eta)}$ strictly increasing in $\eta$. If $\varrho$ happens to take the value $1$, at first glance the context of label shift as in Section 4.2 above reappears. However, as (5) need not hold true under the FJS assumption, in such cases there cannot be label shift. The type of shift modelled instead is called ‘invariant density ratio’ shift (Tasche, 2017), defined as the special case of FJS with $\varrho=1$.
@@ 4.4 Recalibration under assumption of covariate shift with posterior drift (CSPD)
In Sections 4.1, 4.2 and 4.3, the transformation $T$ of (1b) was identified only after a recalibration method had been introduced which was designed under the assumption of one of three types of distribution shift, namely slightly modified covariate shift, label shift, and FJS. In contrast, in this section, $T$ is defined first and then used to characterise the type of distribution shift implied by the recalibration method.
According to Scott (2019), the source distribution $P(X,Y)$ and the target distribution $Q(X,Y)$ are related through covariate shift with posterior drift (CSPD) if there is a strictly increasing transformation $T$ such that (1b) holds true.
As mentioned in Section 3.1, (1b) implies that $\eta_{P}(X)$ is sufficient for $X$ with respect to $Y$ under the target distribution $Q$. Since under CSPD the transformation $T$ is strictly increasing, (1b) also implies that $\eta_{Q}(X)$ and $\eta_{P}(X)$ are strongly comonotonic (Tasche, 2021). As a consequence, Kendall’s $\tau$ and Spearman’s rank correlation both take the maximum value $1$ when applied to $(\eta_{Q}(X),\eta_{P}(X))$. This suggests that CSPD is a strong assumption which might be less often true than one would hope for. Nonetheless, comonotonicity of score and posterior probability is a common assumption in the literature. Chen et al. (2018) call this assumption the rationality assumption.
If labels are available in the test dataset, under the CSPD assumption the transformation $T$ of (1b) can be approximately determined by means of isotonic regression. However, the general assumption for this paper is that there are no label observations for the instances in the test dataset. Therefore, a moment matching approach is applied instead (quasi moment matching, QMM).
To be able to do so, consider parametric CSPD where the transformation $T$ of (1b) is specified as follows through some strictly increasing and continuous distribution function $F$ on the real line and parameters $a,b\in\mathbb{R}$:
@@
This is not a radically new approach but rather a variation of what was called regression-based calibration by Ojeda et al. (2023). Here are two examples for natural choices of $F$ in (8a):
  • Logistic distribution function (inverse logit): $F(x)=\frac{1}{1+\exp(-x)}$, $x\in\mathbb{R}$. The parametric CSPD approach with inverse logit is often called ‘Platt scaling’ in the literature. However, as pointed out by Ojeda et al. (2023), ‘Platt scaling’ sometimes also refers to the specification of $T$ as
@@
    In the following, the term Platt scaling refers to (8b) and logistic CSPD refers to (8a) with inverse logit.
  • Standard normal distribution function (inverse probit): $F(x)=\Phi(x)$, $x\in\mathbb{R}$. This choice of $F$ is referred to as normal CSPD.
@@
In (8c), AUC works like a second moment of the posterior probabilities. But in contrast to its effect on the second moment or the variance, the prior probability of the positive class has no effect on AUC. This makes the assumption of AUC invariance between source and target distributions more plausible (Tasche, 2009).
Remark 4.2
@@
Then the mean of the resulting posterior probability $\eta_{Q}(X)=T_{a,0}(\eta_{P}(X))$ under the target feature distribution $Q(X)$ is computed to obtain an estimate $\widehat{q}$ of $q$:
@@ 4.5 Quasi moment matching based on parametrised receiver operating characteristics
Tasche (2009, Section 5.2) modified an idea of Van der Burgt (2008) by assuming the ROC curve of a real-valued score $S$ to be
@@
for some fixed parameter $c\in\mathbb{R}$, with $\Phi$ denoting the standard normal distribution function. The ROC curve of (10a) emerges when the class-conditional score distributions in a binary classification problem are both univariate normal distributions with equal variances. But ROC curves like in (10a) may also be incurred in circumstances where the class-conditional distributions are not normal (Proposition 5.3 of Tasche, 2009).
Tasche (2009) showed that (10a) implies the following representation for the posterior probability given the score $S$ under the target distribution $Q$:
[…]
@@
Denote by $f$ an unconditional density of $S$ under $Q$ and by $f_{0}$ a density of $F_{0}$. Then the posterior probability Q[Y=1|S=s]Q[Y=1\,|\,S=s] can be represented as (see, e.g., Section 2.4.2 of Tasche, 2017)
@@
In the case of scores $S$ under a discrete or empirical distribution, as described in Appendices A.1 and A.2, (11b) can be treated as a fixed point equation for the probabilities $Q[S=s\,|\,Y=0]=f_{0}(s)$ and be solved by a straightforward fixed-point iteration with intial values $Q[S=s]$, $s\in\mathcal{S}$. The numerical example of Section 5 below suggests that such an iterative approach converges as long as the prior probability $q$ is small and, hence, $f$ and $f_{0}$ are close to each other.
A further issue when making (10b) operational may occur when the class-conditional distribution function $F_{0}$ or any function approximating it takes the value $1$. In particular, this will happen if $F_{0}$ is approximated by an empirical distribution function. Then the term $\Phi^{-1}\bigl(F_{0}(s)\bigr)$ is ill-defined. See Appendix A.2 for a workaround to deal with this issue.
Deploying a discrete distribution $F_{0}$ in (10b) as for instance an empirical approximation, makes it unlikely if not impossible to exactly match a pre-defined $AUC_{S}$ when using (10b). The unavoidable deviation from the $AUC_{S}$ objective in this case can be controlled with the original QMM approach as proposed by Tasche (2009). Define
@@
To distinguish the two recalibration approaches presented in this section, the approach based on (10b) and (10c) is called ROC-based QMM. The approach based on (12a) and (12b) is referred to as 2-parameter QMM.
5 Illustration
The illustrative example presented in this section shows the impact of the recalibration methods and assumptions discussed in Section 4 on the following characteristics of the target distribution: